OpenStudy (anonymous):

Paula had 25 pictures. She wanted to put 5 pictures on each page in her album. How many pages would Paula need? Paula knew she could solve this problem using the number sentence 25 ÷ 5 = . Which is another way Paula could solve this problem? A. Multiply 25 by 5. B. Subtract 5 from 25. Keep subtracting 5 from the answer until she gets to 0. Then count the number of times 5 was subtracted. C. Add 25 and 5. D. Add 5 to 25. Keep adding 5 to the answer a total of five times.

OpenStudy (anonymous):

will fan and medal

OpenStudy (alexandervonhumboldt2):

B

OpenStudy (alexandervonhumboldt2):

medal and fan

OpenStudy (anonymous):

help me @LacedUpCountry @jojo4eva @funfuse2

OpenStudy (lacedupcountry):

Im going to go with B. because 25-5=20 so theres 1 time, 20-5=15 theres 2, 15-5=10 theres 3, 10-5=5 theres 4. and 5-5=0 and thats 5 times.

OpenStudy (lacedupcountry):

Does that makes sense to you?